50% of 90 is the same as 40% of this number

Answer by jsmallt9(3758)

When a problems mentions "[fraction or percent] of [some number]", the "of" translates into a multiplication. And if it says percent, like your problem, use the fraction or decimal equivalent of the percents. (Do not use actual percents in arithmetic!)

So, using 0.50 for 50%, 0.40 for 40% and x for the unknown number, "50% of 90 is the same as 40% of this number" translates into:
0.50 * 90 = 0.40 * x
Now we just solve for x. First we simplify the left side:
45 = 0.40x
Now we divide by 0.40:
45/0.40 = x
which simplifies to:
112.5 = x